# ABACUS: An Electronic Structure Analysis Package for the AI Era

### v1

@@ V.5 Hybrid Funtional
$v(|\mathbf{r}-\mathbf{r}^{\prime}|)\psi^{*}_{j\mathbf{k}^{\prime}}(\mathbf{r}^{\prime})\psi_{i\mathbf{k}}(\mathbf{r}^{\prime}).$
$H^{\text{EXX}}_{U\mu,V\nu}(\mathbf{R}_{UV})=$ $\sum_{KL\mathbf{R}_{KL}}\sum_{kl}\sum_{\mathbf{R}^{\prime}}D_{K\kappa,L\lambda}$ (101)
$(\mathbf{R}_{KL})\langle\phi^{\mathbf{0}}_{U\mu}\phi^{\mathbf{R}^{\prime}}_{K\kappa}|\hat{K}|\phi^{\mathbf{R}_{UV}}_{V\nu}\phi^{\mathbf{R}^{\prime}+\mathbf{R}_{KL}}_{L\lambda}\rangle,$
in which $\hat{K}$ is the exchange operator, the density matrix and the 4-center integrals are defined as
$D_{K\kappa,L\lambda}(\mathbf{R}_{KL})=\frac{1}{N_{\mathbf{k}}}\sum_{j\mathbf{k}}f_{j}c_{j\mathbf{k},K\kappa}c_{j\mathbf{k},L\lambda}^{*}e^{-i\mathbf{\mathbf{k}\cdot\mathbf{R}_{KL}}},$ (102)
@@
$v(\mathbf{r}-\mathbf{r}^{\prime})\phi^{\mathbf{R}_{V}}_{V\nu}(\mathbf{r}^{\prime})\phi^{\mathbf{R}_{L}}_{L\lambda}(\mathbf{r}^{\prime}),$
with $v(\mathbf{r}-\mathbf{r}^{\prime})=1/|\mathbf{r}-\mathbf{r}^{\prime}|$ for the bare Coulomb in PBE0 and HF and $v(\mathbf{r}-\mathbf{r}^{\prime})=\mathrm{erfc(\omega|\mathbf{r}-\mathbf{r}^{\prime}|)}/|\mathbf{r}-\mathbf{r}^{\prime}|$ for short-range EXX in HSE.
Figure 20: Flowchart of the hybrid functional as implemented with numerical atomic orbitals and LRI-based exact exchange calculations in ABACUS.
V.5.1 Local Resolution of Identity
The $O(N^{4})$ scaling EXX effective potential is much more consuming for both time and memory than the local ones. To address this, various types of algorithms have been developed, one of them is the resolution of identity (RI) Feyereisen, Fitzgerald, and Komornicki (1993); Vahtras, Almlöf, and Feyereisen (1993); Ren et al. (2012) or density fitting Whitten (1973) technique, reducing the 4-center integrals to 3- and 2-center ones by expanding the orbital products on a set of auxiliary basis functions (ABFs). Aiming at HDF calculations of large periodic systems, local resolution of identity (LRI)Levchenko et al. (2015); Ihrig et al. (2015); Lin, Ren, and He (2020) can be employed to achieve linear-scaling EXX calculations due to the locality of the atomic orbitals (AOs) and ABFs. If the ABFs are well chosen, the LRI approximation can be adequately accurate for HDF calculations, as benchmark testedLin, Ren, and He (2020). Most of the remaining differences in the obtained cohesive properties and band gaps result from the core–valence interaction treatment (pseudopotentials versus all-electron) rather than the basis set differencesJi et al. (2022).
There are two types of ABFs used in ABACUS: the “on-site” ones to fit the products of orbitals on the same atoms, and the “opt” ones to improve the fitting of different-atom orbital products. Both of them have the same function type as AOs, i.e. a radial function times a spherical harmonics: $P_{A\alpha=\{nlm\}}(\mathbf{r})=g_{nl}(r)Y_{lm}(\hat{\mathbf{r}})$. The radial part of “on-site” ABFs are generated by multiplying the radial part of each AO pair on the same atom: $g_{nl}(r)=f_{n_{1}l_{1}}(r)f_{n_{2}l_{2}}(r),~{}|l_{1}-l_{2}|\leq l\leq l_{1}+l_{2}$, and then orthogonalized and selected by principal component analysis (PCA). The “opt” ABFs are generated in the same way as AO-generationChen, Guo, and He (2010b, 2011b).
Figure 21: The band gap of each system, calculated by ABACUS, along with the reference computed values from other software (QE/FHI-AIMS/VASP) and experimental values, is also presented online ope (2024b).
Due to the locality of the AOs and density matrix, the submatrices involved in calculations are close to zero on many occasions and can be prescreened by a given threshold respectively. There are two ways to distribute the 4-center-integral tasks. One is the greedy algorithm for multi-machine scheduling aiming at load balance, and the other is to distribute the ABFs-located atom pairs according to the K-means clustering result, so that the pairs on the same processor are as close as possible to each other, and the memory consumption in direct proportion to the union of their adjacent tables is minimized Lin, Ren, and He (2021a).
To minimize the copy of tensors between processors, the users are recommended to use thread-level parallelism (OpenMP) within one node and process-level parallelism (MPI) between nodes.
To set up the hybrid functional calculation with LRI-based EXX, the “on-site” ABFs are generated according to the read-in AOs and then orthogonalized and PCA-selected, and the “opt” ABFs are read in if provided. After that, the 2-center radial integral interpolation table between AOs and ABFs are constructed. The above steps are done only once at the beginning of the program.
For each configuration, or in each ion step, the position-dependent tensors $C,V$ are calculated before the SCF iterations, and force and stress are calculated (if needed) after the convergence of SCF iterations. The middle part is composed of at least two SCF iterations: The first one with PBE functional is to get the density matrix to initialize the EXX calculation. After that, there are two different modes to converge the hybrid functional SCF iterations:
Non-separate-loop mode
In the second SCF iteration, the EXX energy and Hamiltonian are updated with the current density matrix in every electronic step. Compared to the separate-loop mode, each electronic step takes longer, but usually, fewer steps are needed to converge the SCF iteration.
Separate-loop mode
In this mode, two layers of loops are used to converge the SCF iteration: The EXX energy and Hamiltonian are updated only at the beginning of the outer loop, which will be fixed and added to the total energy and Hamiltonian in the inner loop (where the density matrix and (semi-)local parts of Hamiltonian are updated). After the inner loop converges, the program checks the outer loop’s convergence and returns to the EXX-update step if it has not converged. The whole workflow of hybrid functional implementation in ABACUS is shown in Fig. 20.
It is also worth mentioning that the LRI-based 4-center integration step has been individually packaged into a library named LibRIlib (2024a), which can be integrated into other DFT or beyond-DFT (such as RPA and GW) codeslib (2024b) to speed up the calculation of the two-electron Coulomb repulsion integrals.
Figure 22: Band gaps of Cs2AgInCl6 with a lattice constant of 10.481 Åpredicted by HSE(0.4), HSE, SCAN, DFT+U,PBE,and PBEsol as a function of Ag-Cl bond length. (Adapted with permission from Phys. Rev. Research 6, 033172 (2024). Copyright 2024 American Physical Society.)
V.5.2 Results
[…]
Figure 23: Time consumption per iteration of three main parts in ABACUS SCF calculations using hybrid functional. Here HFX represents the HFX matrix construction, and Rest means the rest of time. (Adapted with permission from J. Chem. Theory Comput. 2021, 17, 1, 222–239. Copyright 2021 American Chemical Society.)
@@ V.5.2 Results
Generally speaking, hybrid functional calculations can only process systems of limited size due to the large computational demand of building the exact exchange Hamiltonian. By combining NAOs basis and LRI techniques, Fig. 21 shows the time consumption of this part increases almost linearly with the system size, enabling ABACUS to perform hybrid functional calculations for systems with thousands of atoms. In 2024, Tang et al. used hybrid functional data from ABACUS and trained a deep equivariant neural network approach for efficient hybrid density functional calculations Tang et al. (2024). Taking twisted bilayer graphene (TBG) as an example in Fig. 24, they performed the HSE band structure calculations of (17, 16) TBG (twist angle $\theta\approx 2.0046^{\circ}$, 3,268 atoms/cell)Tang et al. (2024).
@@
We alo perform HSE calculations to obtain the band gaps of several semiconductors and compare them with the results from other DFT packages (QE/VASP/FHI-AIMS) and the experimental values Lin, Ren, and He (2020); Cui et al. (2018). In these tests, ABACUS and QE employ the SG15-type norm-conserving PBE pseudopotentials (except for In is PSlibrary norm-conserving), VASP employs the PBE PAW potentials, whereas FHI-AMIS performs all-electron calculations Delley (1990). Fig. 21 shows that the band gaps obtained by ABACUS HSE calculations are consistent with the results from other DFT packages, demonstrating satisfactory results when compared to the experimental values.
@@
Figure 24: The PBE and HSE06 band structures of twisted bilayer graphene with $\theta\approx 2.0046^{\circ}$, 3268 atoms/cell computed by ABACUS. (Adapted with permission from Nat Commun 15, 8815 (2024). Copyright 2024 Springer Nature.)
V.6 The DeePKS Method
@@
Density functional theory has been widely used to calculate the electronic structural properties of molecules and condensed systems. The exchange-correlation functional plays a crucial role in determining the accuracy of DFT. However, when selecting different levels of XC functionals, there is a trade-off between accuracy and efficiency. A lower-level XC functional in the Jacob ladderPerdew (2001) usually has higher computational efficiency but lower accuracy, while a higher-level one has the opposite. With the rapid advancement of extensive computational resources and advanced algorithms, Artificial Intelligence assisted methods have the potential to help solve this problemKulik et al. (2022); Huang, Von Rudorff, and Von Lilienfeld (2023).
@@
Proposed in 2020, DeeP Kohn-Sham method (DeePKS)Chen et al. (2021) employs a computationally efficient neural network-based functional model to represent the difference between a lower-level XC functional and a higher-level XC functional. The resulting model maintains translational, rotational, and permutational invariance and can be used in self-consistent field calculations. Together with DeePKS-kitChen et al. (2022a) software, ABACUS supports the iterative training of DeePKS modelsLi et al. (2022a). It can be trained in molecular or periodic systems and supports training for properties such as energy, force, stress, and bandgap. Based on the DeePKS model, ABACUS can achieve accuracy similar to that of the high-level XC functional on these specified properties, while the computational efficiency is similar to that of the low-level XC functional.
@@
We divide the energy functional into two parts:
@@
where $\{\psi_{i}\}$ are single-particle orbitals, and $E_{\delta}$ is constructed as a neural network model with parameters $\omega$.
The model input is constructed based on the projected density matrix:
@@
Here $\rho_{\mu\nu}$ represents the density matrix, and $|\alpha\rangle$ is a set of localized orbitals centered on atoms, identified by atomic index $I$, and quantum numbers $nlm$. The atomic-centered basis functions $|\alpha\rangle$ ensure the translational invariance. To maintain rotational invariance, we proceed to extract the eigenvalues of projected density matrix blocks with the same indices $I$, $n$, and $l$, resulting in a set of descriptors:
@@
The descriptors are grouped into vectors based on the atomic index $I$, and $E_{\delta}$ is calculated as the sum of atomic contributions:
@@
which guarantees the invariance under permutations. $F_{\mathrm{NN}}$ is the deep neural network.
Here, with the expression of the energy functional in mind, the Hamiltonian operator can be written as :
$H=H_{\text{baseline}}+\hat{V}^{\delta}.$ (108)
The matrix elements of the correction potential are as follows:
$\hat{V}_{\mu\nu}^{\delta}$ $=\frac{\partial E_{\delta}}{\partial\rho_{\mu\nu}}$
$=\sum_{Inlmm^{\prime}}\frac{\partial E_{\delta}}{\partial D_{nlmm^{\prime}}}\frac{\partial D_{nlmm^{\prime}}^{I}}{\partial\rho_{\mu\nu}}$
$=\sum_{Inlmm^{\prime}}\langle\chi_{\mu}|\alpha_{nlm}^{I}\rangle\frac{\partial E_{\delta}}{\partial D_{nlmm^{\prime}}^{I}}\langle\alpha_{nlm^{\prime}}^{I}|\chi_{\nu}\rangle.$ (109)
A trained DeePKS model provides the chance to perform KS-DFT calculations with high computational accuracy and efficiency at the same time. It can be applied to precisely predict various properties for a family of systems and potentially facilitate the discovery and design of novel materials.
For instance, Ou et al.Ou et al. (2023) established a general DeePKS model that can be utilized for a plethora of halide perovskites, including different combinations of ABX3 (A=FA, MA, Cs; B=Sn, Pb; X=Cl, Br, I), the organic-inorganic hybrid alternatives and the Ruddlesden-Popper (RP) perovskites. Halide perovskites (ABX3, X=halogen anion) have shown great promise as a cost-effective alternative to current commercial photovoltaic technologies. One key benefit of halide perovskites is adjusting the absorption edge wavelength (band gap) by changing the ratio of different halide ions. Designing effective photovoltaic systems requires a precise yet efficient description of the electronic structure of halide perovskites.
They built an extensive DeePKS model upon 460 configurations spanning seven types of halide perovskites, with HSE06 accuracy and satisfactory predictions for the band gap. Based on an iterative training process with ABACUS and DeePKS-kit, they showed that the resulting DeePKS model can accurately replicate forces, stress, band gaps, and density of states (DOS) near the Fermi energy for all types of halide perovskites, including RP structures, and hybrid compositions, when compared to HSE06. For example, they show the band gaps predicted by DeePKS and PBE, with respect to those by HSE06, over 30 tested systems. As depicted in Fig. 25, The DeePKS model demonstrated precise predictions of band gaps for all perovskites examined, closely aligning with the HSE06 findings and yielding an average absolute error (MAE) of 0.0350 eV. PBE significantly underestimated the band gap values, with a large MAE of 0.5222 eV.
@@
In addition to providing an accurate yet efficient description of the electronic structure, DeePKS also serves as a “bridge” between expensive quantum mechanical (QM) models and machine learning (ML)-based potentials. While the ML-based potentials such as the Deep Potential Molecular Dynamics (DeePMD)Zhang et al. (2018a); Wang et al. (2018) have emerged as powerful tools for mitigating the high computational costs associated with ab initio molecular dynamics (AIMD), training these potentials demands a significant number of QM-labeled frames. DeePKS offers a solution to further save the computational cost by reducing the required QM-labeled frames, owing to its significantly better transferability as compared to DeePMD.
Li et al. Li et al. (2022a) examined DeePKS and DeePMD’s training curves with respect to training samples in systems with 64 water molecules at the accuracy of hybrid functional SCAN0. Evidently, the DeePKS model outperforms the DeePMD model with fewer frames as shown in Fig. 26. Additionally, the DeePKS model has a smaller generalization gap than the DeePMD model. Also, the DeePKS model can accurately replicate the SCAN0 SCF result for 64 water molecules in just fifteen minutes, significantly improving from the original time of over a dayLi et al. (2022a).
Figure 26: Learning curves for energy (upper panel) and force (lower panel) given by DeePMD (orange) and DeePKS (blue) with respect to the number of training frames.Li et al. (2022a) Dashed line with squares indicates train set error; solid line with diamonds indicates test set error. (Adapted with permission from J. Phys. Chem. A 2022, 126, 49, 9154–9164. Copyright 2022 American Chemical Society.)
@@
The work Li et al. (2022a) also showed that excellent agreement can be achieved between the ABACUS-DeePKS-DeePMD results and the ones from SCAN or SCAN0-based DeePMD simulations. By labeling fewer than 200 frames in the training set with hybrid meta-GGA SCAN0 or meta-GGA SCAN functionals, they have shown that the GGA-based DeePKS model can effectively reproduce the energies and forces for pure and salt water systems, with considerable time savings. SCF calculations with trained DeePKS models were carried out and utilized as labels for DeePMD training. Liquid water’s structural properties, such as radial distribution function (RDF), bulk density, H-bonds, and dynamic properties like diffusion coefficient, were found to be excellently matched with those obtained by the SCAN0 AIMD and DeePMD methods. For example, for systems consisting of 64 water molecules, they trained a DeePKS model with 180 training samples at the accuracy of SCAN0. Then, they used it to quickly develop 1000 data for training DeePMD potential and performed MD of 512 water molecules. As shown in Fig. 27, various radial distribution functions (RDFs) derived from DeePKS-DeePMD simulations exhibit remarkable consistency with both the SCAN0-AIMD and SCAN0-DeePMD results, including a marked decrease in overstructured peaks.
@@
Figure 27: Radial distribution functions (RDFs) (a) $g_{\textrm{OO}}(r)$, (b) $g_{\textrm{OH}}(r)$, and (c) bond angle distribution $P_{\textrm{OO}}(\theta)$ given by DeePKS-DeePMD (blue dotted line), SCAN0-AIMD (black solid line), SCAN0-DeePMD (orange dashed line), and PBE-AIMD (gray dotted-dashed line).(Adapted with permission from J. Phys. Chem. A 2022, 126, 49, 9154–9164. Copyright 2022 American Chemical Society.)
@@
Zhang et al. Zhang et al. (2024b) investigated the tautomeric equilibria of glycine in water with the DeePMD model at the accuracy of M06-2XZhao and Truhlar (2008). To avoid expensive computational cost for directly generating dataset at M06-2X level, they utilized the DeePKS method as implemented in ABACUS and DeePKS-kit, together with on-the-fly probability enhanced sampling (OPES)Invernizzi and Parrinello (2020) method, to construct the dataset for training the DeePMD model. With this DeePMD potential and OPES, they performed MD and got a converged free energy surface (FES). They observed that glycine can undergo tautomerism, transitioning between its neutral and zwitterionic forms through intramolecular and intermolecular proton transfers. In Zhang and Xu (2024), Zhang et al. adopted a similar strategy to study the propensity of water self-ions at air(oil)-ater interface. They found that the trained DeePKS model can decrease calculation time by about nine times compared to regular M06-2X calculations. With the resulting DeePMD model for an efficient MD process, they demonstrated the stable ionic double-layer distribution near the interface for both air-water and oil-water interface systems.
V.7 DFT+U

### v3

@@ IV.4 Hybrid Functional
Figure 19: Band gaps of Cs2AgInCl6 with a lattice constant of 10.481 Å, as computed by HSE (0.4), HSE, SCAN, DFT+U, PBE ,and PBEsol as a function of Ag-Cl bond length. (Adapted with permission from Phys. Rev. Research 6, 033172 (2024). Copyright 2024 American Physical Society.)
Figure 20: Band gaps of 21 different semiconductors, calculated by ABACUS+LibRI, along with the reference computed values from other software (QE/FHI-AIMS/VASP) Lin, Ren, and He (2020) and experimental values Cui et al. (2018).
The computational efficiency is substantially enhanced by exploiting the inherent locality of both AOs and the density matrix, which enables effective prescreening of negligible submatrices below a specified threshold. For parallel computation of four-center integrals, we have implemented two complementary distribution strategies: a load-balancing greedy algorithm for multi-processor scheduling, and a spatially optimized scheme based on K-means clustering of ABF atom pairs. The latter minimizes both inter-processor communication overhead and memory requirements by ensuring geometrically proximate pairs are assigned to the same processor, with memory consumption scaling linearly with the union of their neighbor lists Lin, Ren, and He (2021a).
Notably, the LRI-based four-center integration has been modularized into the standalone LibRI library lib (2024a), which provides a portable and efficient implementation that can be readily integrated into various electronic structure codes. This includes both conventional DFT packages and beyond-DFT methods such as the random phase approximation (RPA) and GW calculations lib (2024b), offering substantial acceleration for two-electron Coulomb repulsion integral evaluations across different theoretical frameworks.
In 2025, Lin et al. developed a linear-scaling algorithm for exact-exchange forces and stresses in hybrid DFT, applicable to both molecules and periodic systems. Leveraging numerical atomic orbitals and localized resolution-of-identity, the massively parallel implementation enables structural relaxation of thousand-atom systems in ABACUS Lin et al. (2025).
@@
Generally speaking, hybrid functional calculations can only process systems of limited size due to the large computational demand of building the exact exchange Hamiltonian. By combining NAOs basis and LRI techniques, Fig. 20 shows the time consumption of this part increases almost linearly with the system size, enabling ABACUS to perform hybrid functional calculations for large systems with thousands of atoms. In 2024, Tang et al. used hybrid functional data from ABACUS and trained a deep equivariant neural network approach for efficient hybrid density functional calculations Tang et al. (2024). Taking twisted bilayer graphene (TBG) as an example in Fig. 21, they performed the HSE band structure calculations of (17, 16) TBG (twist angle $\theta\approx 2.0046^{\circ}$, 3,268 atoms/cell)Tang et al. (2024).
@@
We also performed HSE calculations to obtain the band gaps of several semiconductors and compare them with the results from other DFT packages (QE/VASP/FHI-AIMS) and the experimental values Lin, Ren, and He (2020); Cui et al. (2018). In these tests, ABACUS and QE employ the SG15-type norm-conserving PBE pseudopotentials (except for In is PSlibrary norm-conserving), VASP employs the PBE PAW potentials, whereas FHI-AMIS performs all-electron calculations Delley (1990). Fig. 20 shows that the band gaps obtained by ABACUS HSE calculations are consistent with the results from other DFT packages, demonstrating satisfactory results when compared to the experimental values.
@@
Figure 21: Band structures of twisted bilayer graphene with $\theta\approx 2.0046^{\circ}$, 3268 atoms/cell computed by ABACUS using the PBE and HSE06 functionals. (Adapted with permission from Nat Commun 15, 8815 (2024). Copyright 2024 Springer Nature.)
IV.5 DeePKS
@@
The exchange-correlation functional is critical to determining the accuracy of DFT. However, when different levels of XC functionals are selected, there exists a trade-off between accuracy and efficiency. Lower-level XC functionals on Jacob’s ladder Perdew (2001) typically offer higher computational efficiency but lower accuracy, whereas higher-level ones exhibit the opposite trend. With the rapid progress in increasing computational resources and advanced algorithms, AI-assisted methods hold the potential to mitigate this issue Kulik et al. (2022); Huang, Von Rudorff, and Von Lilienfeld (2023).
@@
Proposed in 2020, DeeP Kohn-Sham method (DeePKS)Chen et al. (2021) employs a computationally efficient neural network-based functional model to represent the energy and force difference between a lower-level XC functional and a higher-level XC functional. The resulting model maintains translational, rotational, and permutational invariance and can be used in self-consistent field calculations. Together with DeePKS-kit Chen et al. (2022a) software, ABACUS supports the iterative training of DeePKS models Li et al. (2022a). Specifically, DeePKS models can be trained in molecular or periodic systems to learn properties such as energy, force, stress, and band gaps. Based on the DeePKS model, ABACUS can achieve accuracy similar to that of the high-level XC functional on these specified properties, while the computational efficiency is similar to that of the low-level XC functional.
@@
In the DeePKS method, we divide the total energy functional into two parts
@@
where $\{\psi_{i}\}$ are single-particle orbitals to yield the baseline energy $E_{\text{baseline}}$, and $E_{\delta}$ is constructed as a neural network model with parameters $\omega$. The model input is constructed based on the projected density matrix for each atom $I$, which takes the form
@@
Here $\rho_{\mu\nu}$ represents the density matrix, and $|\alpha_{nlm}^{I}\rangle$ is a set of localized orbitals centered on atoms, identified by atomic index $I$, and quantum numbers nlmnlm. The atomic-centered basis functions $|\alpha_{nlm}^{I}\rangle$ ensure the translational invariance. To maintain rotational invariance, we proceed to extract the eigenvalues of projected density matrix blocks with the same indices $I$, $n$, and $l$, resulting in a set of descriptors that are obtained as eigenvalues of the projected density matrix
@@
The descriptors are then grouped into vectors based on the atomic index $I$, and $E_{\delta}$ is calculated as the sum of atomic contributions as
@@
where $F_{\mathrm{NN}}$ is the deep neural network. Next, the Hamiltonian operator can be written as
$H=H_{\text{baseline}}+\hat{V}^{\delta},$ (96)
where the $H_{\text{baseline}}$ term is the baseline Hamiltonian, and the correction potential can be evaluated via
$\hat{V}_{\mu\nu}^{\delta}=\frac{\partial E_{\delta}}{\partial\rho_{\mu\nu}}.$ (97)
Finally, the corresponding force and stress terms can also be obtained for the correction term. Li et al. (2022a)
Here we list a few applications using the DeePKS method. For instance, halide perovskites (ABX3, X=halogen anion) have shown great promise as a cost-effective alternative to current commercial photovoltaic technologies. Designing effective photovoltaic systems requires a precise yet efficient description of the electronic structure of halide perovskites. One key benefit of halide perovskites is adjusting the absorption edge wavelength (band gap) by changing the ratio of different halide ions. Ou et al. Ou et al. (2023) constructed a general DeePKS model that can be utilized for halide perovskites, including different combinations of ABX3 (A=FA, MA, Cs; B=Sn, Pb; X=Cl, Br, I), the organic-inorganic hybrid alternatives, and the Ruddlesden-Popper (RP) perovskites. In detail, the authors built an extensive DeePKS model upon 460 configurations spanning seven types of halide perovskites, with HSE06 accuracy and satisfactory predictions for the band gap. Based on an iterative training process with ABACUS and DeePKS-kit, they showed that the resulting DeePKS model can accurately replicate forces, stress, band gaps, and density of states (DOS) near the Fermi energy for all types of halide perovskites, including RP structures, and hybrid compositions, when compared to HSE06. They further show the band gaps predicted by DeePKS and PBE, with respect to those by HSE06, over 30 tested systems. As depicted in Fig. 22, The DeePKS model demonstrated precise predictions of band gaps for all perovskites examined, closely aligning with the HSE06 findings and yielding an average absolute error (MAE) of 0.0350 eV. On the other hand, PBE significantly underestimated the band gap values, with a large MAE of 0.5222 eV.
@@
In addition to providing an accurate yet efficient description of the electronic structure, DeePKS also serves as a “bridge” between expensive quantum mechanical (QM) models and ML-based potentials. While the ML-based potentials such as the Deep Potential Molecular Dynamics (DeePMD)Zhang et al. (2018a); Wang et al. (2018) have emerged as powerful tools for mitigating the high computational costs associated with ab initio molecular dynamics (AIMD), training these potentials demands a significant number of QM-labeled frames. DeePKS offers a solution to further save the computational cost by reducing the required QM-labeled frames, owing to its significantly better transferability as compared to DeePMD. Li et al. Li et al. (2022a) examined DeePKS and DeePMD’s training curves with respect to the number of training samples in systems with 64 water molecules at the accuracy of hybrid functional SCAN0. The DeePKS model outperforms the DeePMD model with fewer frames shown in Fig. 23. Additionally, the DeePKS model has a smaller generalization gap than the DeePMD model. The work Li et al. (2022a) also showed that excellent agreement can be achieved between the ABACUS-DeePKS-DeePMD results and the ones from SCAN or SCAN0-based DeePMD simulations. By labeling fewer than 200 frames in the training set with SCAN0 or SCAN functionals, the GGA-based DeePKS model can effectively reproduce the energies and forces for pure and salt water systems, with considerable time savings. SCF calculations with trained DeePKS models were carried out and utilized as labels for DeePMD training. Structural properties of liquid water, such as radial distribution function (RDF), bulk density, hydrogen (H) bonds, and dynamic properties like diffusion coefficient, were found to be excellently matched with those obtained by the SCAN0 AIMD and DeePMD methods. For example, for systems consisting of 64 water molecules, they trained a DeePKS model with 180 training samples at the accuracy of SCAN0. Then, the authors used it to efficiently generate 1000 data to train DeePMD potential and performed MD simulations of 512 water molecules. As shown in Fig. 24, various RDFs derived from DeePKS-DeePMD simulations match well with both SCAN0-AIMD and SCAN0-DeePMD results, including a marked decrease in overstructured peaks.
Figure 23: Learning curves for energy (upper panel) and force (lower panel) given by DeePMD (orange) and DeePKS (blue) with respect to the number of training frames. Li et al. (2022a) Dashed line with squares indicates train set error, while solid line with diamonds indicates test set error. (Adapted with permission from J. Phys. Chem. A 2022, 126, 49, 9154–9164. Copyright 2022 American Chemical Society.)
@@
Figure 24: Radial distribution functions (RDFs) of (a) $g_{\textrm{OO}}(r)$, (b) $g_{\textrm{OH}}(r)$, and (c) bond angle distribution $P_{\textrm{OO}}(\theta)$ given by DeePKS-DeePMD (blue dotted line), SCAN0-AIMD (black solid line), SCAN0-DeePMD (orange dashed line), and PBE-AIMD (gray dotted-dashed line).(Adapted with permission from J. Phys. Chem. A 2022, 126, 49, 9154–9164. Copyright 2022 American Chemical Society.)
@@
Zhang et al. Zhang et al. (2024b) investigated the tautomeric equilibria of glycine in water with the DeePMD model at the accuracy of M06-2XZhao and Truhlar (2008). To avoid expensive computational cost for directly generating dataset at M06-2X level, they utilized the DeePKS method as implemented in ABACUS and DeePKS-kit, together with on-the-fly probability enhanced sampling (OPES)Invernizzi and Parrinello (2020) method, to construct the dataset for training the DeePMD model. With this DeePMD potential and OPES, they performed MD simulations and obtained a converged free energy surface (FES). They observed that glycine can undergo tautomerism, transitioning between its neutral and zwitterionic forms through intramolecular and intermolecular proton transfers. In Zhang and Xu (2025), Zhang et al. adopted a similar strategy to study the propensity of water self-ions at air(oil)-water interface. They found that the trained DeePKS model can decrease calculation time by about nine times compared to regular M06-2X calculations. With the resulting DeePMD model for an efficient MD process, they demonstrated the stable ionic double-layer distribution near the interface for both air-water and oil-water interface systems.
Liang et al. Liang, Liu, and Chen developed an enhanced version of the DeePKS approach, known as DeePKS-ES (electronic structure), to further improve its prediction of electronic structure properties. The DeePKS-ES method modifies the loss function by integrating the Hamiltonian matrix along with its eigenvalues (energy levels) and eigenvectors (wave function coefficients). This adjustment allows DeePKS-ES to reliably predict key electronic structure characteristics like band gaps, density of states (DOS), and Hamiltonian matrices, in addition to total energy and atomic forces. When tested on various water systems, including monomers, clusters, and liquid phases, DeePKS-ES achieves a level of accuracy similar to that of the hybrid functional HSE06 while maintaining the computational efficiency of the generalized gradient approximation PBE. Particularly, the universal DeePKS-ES model shows strong adaptability across different water systems, connecting quantum-mechanical precision with scalable computation.
IV.6 DFT+U